OpenStudy (anonymous):

sale price per item is $15, cost of producing item is $55 and labor per item is $4. What is the equation to determine revenue, total cost and how many items must be sold to break even?

OpenStudy (anonymous):

Mollyann, I am mathematically illiterate....please explain in detail?

OpenStudy (anonymous):

The cost of production isn't per-item, is it?

OpenStudy (anonymous):

yes it is.

OpenStudy (anonymous):

If it costs 55$ to produce and sells for 15$ then it is impossible to make a profit. This is a losing proposition.

OpenStudy (anonymous):

I'm going to continue with the assumption that the 55$ is a one-time cost - otherwise, no business would actually do this... Anyway, the basic equations are: Revenue = Price × number of items Profit = Revenue - Total Cost Total Cost = one-time (I hope) Production Cost + Labor Cost Break-Even is when Profit = 0 or Revenue = Total Cost

OpenStudy (anonymous):

If you let x = number of items sold, then Total Cost = 55+4x, and Revenue = 15x Can you figure it out from here?

OpenStudy (anonymous):

No, as I am mathematically illiterate when it comes to algebra. I need all of the help that I can get when it comes to college algebra.

OpenStudy (anonymous):

Hmm, well there isn't much algebra to do here. It is more a matter of writing the expressions correctly. You'll have to solve a simple equation for the finding the break-even point, but not much else.

OpenStudy (anonymous):

Since break-even occurs when Total Cost = Revenue, and Total Cost = 55+4x, and Revenue = 15x Your equation to solve is 55+4x=15x. Can you handle that?

OpenStudy (anonymous):

Got it, thanks Cliff, you have been the greatest help!!!!

OpenStudy (anonymous):

If you can find it in a library or book store, I recommend the book, Forgotten Algebra; it covers a lot of stuff like this.